Use Substitutionto solve each euqations, If the system does not have exactly one solution, state whethter it has no solution or infinitely many solutions.
X-5Y=36 2X+Y=-16

Answer by stanbon(75887)

Use Substitutionto solve each euqations, If the system does not have exactly one solution, state whethter it has no solution or infinitely many solutions.
X-5Y=36
2X+Y=-16
---------
Rearrange:
x = 5y+36
Substitute:
2(5y+36) + y = -16
10y + 72 + y = -16
11y = - 88
y = -8
---------------
Substitute into x = 5y+16
x = 5*-8+16
x = -40+16
x = -24
===============
One solution: (-24,-8)
===============
Cheers,
Stan H.